# Totally elliptic surface group representations

### v1

@@ Abstract.
We characterize all totally elliptic surface group representations into $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ by showing that they are either representations into a compact subgroup or Deroin–Tholozan representations.
@@ 1. Introduction
1.1. Main result
Surface group representations are homomorphisms from the fundamental group of a surface $\Sigma$ into a Lie group $G$. Their deformation spaces are called character varieties. A character variety is not necessarily a connected space. Dynamical behaviour and qualitative properties of representations may vary considerably between components. At one end of the spectrum are the higher Teichmüller spaces which are a generalization of the classical and extensively studied Teichmüller space of $\Sigma$. Higher Teichmüller spaces, as defined in [Wie18], consist of discrete and faithful representations and tend to behave like moduli spaces of geometric structures on $\Sigma$, although a precise correspondence is still unavailable. This article aims at drawing the reader’s attention towards the other end of the spectrum where connected components are mostly made of non-discrete representations.
We’ll focus on a particular prototype of representations that are antipodal to discrete and faithful ones: the so-called totally elliptic representations.
@@ 1.1. Main result
A surface group representation $\pi_{1}\Sigma\to G$ is totally elliptic if every simple closed curve on $\Sigma$ is mapped to an elliptic element of $G$. By an elliptic element of $G$, we mean an element of a compact subgroup of $G$ [^1] Those elements are sometimes called compact in the literature.
The surface $\Sigma$ will always be taken to be oriented and connected, and can be punctured. Recall that a simple closed curve on $\Sigma$ is the free isotopy class of a closed loop on $\Sigma$ that doesn’t self-intersect and isn’t homotopically trivial. A simple closed curve can be homotopic to a puncture, in which case we call it peripheral. By definition, the images of every peripheral curve by a totally elliptic representation are elliptic. We want to avoid the situation where peripheral curves are mapped to the identity and we say that a representation is reduced when this isn’t the case (Definition 2.7).
The first instances of totally elliptic representations, though not the most interesting, are those into a compact subgroup of $G$. Interestingly, however, by requiring only simple closed curves to be mapped to elliptic elements (as opposed to all curves), one can construct Zariski dense examples of totally elliptic representations into non-compact Lie groups, such as every Hermitian Lie group (Section 1.3), whose deformations remain totally elliptic.
The initial case to investigate is when $G$ is $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$. In that case, the first totally elliptic representations with Zariski dense image were discovered by Benedetto–Goldman [BG99], and later generalized by Deroin–Tholozan [DT19]. Those representations were originally named supra-maximal in [DT19], but we’ll refer to them as DT representations (Definition 2.10), to avoid any confusion with maximal representations from higher Teichmüller theory. DT representations only exist when $\Sigma$ is a genus-0 surface with at least three punctures. As Deroin–Tholozan observed, their deformation spaces are compact components inside the corresponding relative character varieties and made exclusively of totally elliptic representations (Theorem 2.9). The question was raised whether they’re the only totally elliptic representations into $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, besides representations into compact subgroups. We bring a positive answer to that question.
@@
Let $\Sigma$ be a connected and oriented surface with genus $g\geq 0$ and $n\geq 0$ punctures. Reduced totally elliptic representations $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ can only be of two kinds:
  • The image of $\rho$ is contained in a compact subgroup of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$. In that case, we say that $\rho$ is an orthogonal representation.
  • If $\rho$ is not orthogonal, then $g=0$, $n\geq 3$, and $\rho$ is a DT representation.
We prove Theorem A in Section 2.
@@ 1.2. Motivations
Characterizing a certain type of surface group representations (such as totally elliptic ones) by the images of simple closed curves may seem baroque at first. Similar problems, however, have sparked substantial interest in recent decades. Bowditch famously raised the question of whether discrete and faithful representations of a closed surface group $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$—also known as Fuchsian representations—can be characterized by the property of mapping every simple closed curve on $\Sigma$ to a hyperbolic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ (an element whose trace is larger than $2$ in absolute value) [Bow98, Question C]. If true, this would be a finer characterization of Fuchsian representations which are known to be the only representations that map every non-trivial element of $\pi_{1}\Sigma$ to hyperbolic elements of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$. Bowditch’s question has been answered positively for surfaces of genus $2$ by Marché–Wolff [MW16, MW19] and remains open in higher genera.
@@
Of course, not every representation is Fuchsian. Goldman proved that the character variety of representations $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, where $\Sigma$ is a closed surface of genus $g\geq 2$, has $34g-$ 3 connected components [Gol88]. Two of the components are made of conjugacy classes of Fuchsian representations, forming two copies of the Teichmüller space of $\Sigma$; the other $54g-$ 5 components are known as intermediate components and they don’t contain any Fuchsian representation. Bowditch’s question is equivalent to asking whether every representation whose conjugacy class lies in an intermediate component maps at least one simple closed curve on $\Sigma$ to a non-hyperbolic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$. Marché–Wolff also proved that a positive answer to Bowditch’s question would imply that the mapping class group action on every intermediate component is ergodic [MW16, Theorem 1.6]. The latter statement is known as Goldman’s ergodicity conjecture [Gol06, Conjecture 3.1] and is open, like Bowditch’s question, for surfaces of genus at least $3$.
@@
Goldman also conjectured that almost every non-Fuchsian closed surface group representation $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ should be the holonomy of a branched hyperbolic structure on $\Sigma$ which is the data of hyperbolic metric of $\Sigma$ with several conical singularities of angle $2\pi k$ with $k\geq 2$. To be precise, according to Goldman, every dense representation whose conjugacy class lies in an intermediate component should be the holonomy of a branched hyperbolic structure [Gol06, Conjecture 3.9]. Mathews proved some partial implication from a positive answer to Bowditch’s question to Goldman’s geometrization conjecture for intermediate components of co-Euler class $1$ [Mat12]. Observe as well that Goldman’s ergodicity conjecture implies a “full measure version” of the geometrization conjecture, as well as a positive answer to Bowditch’s question on a full measure subset of intermediate components.
@@
Goldman’s two conjectures and Bowditch’s question, and the interplay among them, constitute a whole program to study intermediate components which we refer to as the Bowditch–Goldman program.
@@
Totally elliptic representations, and more particularly DT representations, naturally arise as a prototype of surface group representations that fit perfectly in the Bowditch–Goldman program. By definition, they map every simple closed curve to an elliptic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, and therefore satisfy the analogue of Bowditch’s condition on intermediate representations in a strong sense. The ergodicity of the pure mapping class group action on the character variety of DT representations has been established in [Mar22]. As explained in [DT19, Section 4] and further detailed in [FM23], DT representations are also holonomies of hyperbolic cone structures with branch point singularities, thereby satisfying a precise version of the geometrization conjecture.
1.3. Perspectives away from $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$
@@
This note is also an opportunity to survey some recent developments about totally elliptic representations into Lie groups $G$ different than $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$. Several families of compact connected components of totally elliptic representations were identified when $G$ is a Hermitian Lie group. They’re all genus-0 surface group representations. The first examples were found by Tholozan–Toulisse when $G$ is the Lie group $\mathrm{SU}(p,q)$, as well as when $G$ is $\mathrm{Sp}(2n,\operatorname{\mathbb{R}})$ or $\mathrm{SO}^{\ast}(2n)$ [TT21]. Analogous compact components have been identified by Feng–Zhang when $G$ is the identity component of $\mathrm{SO}(2,n)$ [FZ23]. Their methods studies the topology of some moduli spaces of parabolic Higgs bundles which relate to relative character varieties via the non-abelian Hodge correspondence. In particular, they proved that for every representation $\rho\colon\pi_{1}\Sigma\to G$ whose conjugacy class belongs to one of the compact components they identified, and for every punctured Riemann sphere structure on $\Sigma$, there exists a $\rho$-equivariant holomorphic map from the universal cover $\widetilde{\Sigma}$ to the symmetric space $G/K$. The same property for DT representations was proven by Mondello [Mon16]. Surface group representations satisfying this condition are sometimes called universal variations of Hodge structures.
@@
Unlike DT representations which are all Zariski dense, it’s not clear yet whether every compact component of totally elliptic representations into Hermitian Lie groups other than $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ contains a Zariski dense representation (almost all of them do though, by the preponderance of Zariski dense representations in any character variety). The topology of those compact components is also yet to be determined, whereas DT components are known to be homeomorphic (even symplectomorphic) to complex projective spaces [DT19, Theorem 4].
@@
There are however more explicit examples of compact components of totally elliptic representations discovered by Marc-Zwecker [MZ24]. Marc-Zwecker’s examples are made of representations of a pair of pants into $\mathrm{SU}(2,1)$ where the peripheral curves are mapped to products of reflections through three complex geodesic lines in the complex hyperbolic space of real dimension four. The three complex geodesic lines are the mirrors and they’re arranged as a triangle. The resulting components are homeomorphic to 2-dimensional spheres. The geometric nature of the representations in Marc-Zwecker’s components may lead to other compact components for spheres with a higher number of punctures obtained by “chaining” the triangles of mirrors in the spirit of [Mar24].
@@
A natural question to ask is whether the class of Lie groups for which there exist substantial connected components entirely made of totally elliptic genus-0 surface group representations is larger than the class of Hermitian Lie groups.
Question 1.
@@
Are there semisimple real Lie groups other than Hermitian Lie groups for which there exist compact connected components (of maximal expected dimension) inside relative character varieties of genus-0 surface group representations made of totally elliptic representations and containing Zariski dense representations?
There are reasons to believe that the key property is the presence of a nonempty open subset of elliptic elements in the target Lie group. This property is shared by all Hermitian Lie groups, but not only; for instance, Sugiura proved that $\mathrm{SO}(2p,q)$ has an open subset of elliptic elements [Sug59, Theorem 8], even though it’s not Hermitian if $p\geq 2$ and $q\geq 3$. In contrast, the subset of elliptic elements in $\mathrm{SL}_{n}\operatorname{\mathbb{R}}$ has empty interior for every $n\geq 3$. Furthermore, Kabenyuk proved that admitting a nonempty open subset of elliptic elements is equivalent to having a compact Cartan subgroup [Kab90].
@@
One should nevertheless be aware that “large” Lie groups, unlike $\mathrm{SL}_{2}\operatorname{\mathbb{C}}$, may admit subgroups $\Gamma$ containing only elliptic elements but such that $\Gamma$ is not contained into a compact subgroup. Bass gives an example of such a $\Gamma$ which is isomorphic to a rank-2 free subgroup of $\mathrm{SL}_{3}\operatorname{\mathbb{C}}$ [Bas80, Counterexample 1.10]. Bass’ example is not Zariski dense and, in fact, it couldn’t be as explained for instance in [Pra94]. However, when $\operatorname{\mathbb{C}}$ is replaced by a non-Archimedean valued field $\mathbb{F}$, then the situation is similar to the case of $\mathrm{SL}_{2}\operatorname{\mathbb{C}}$ as Parreau proved that a finitely generated subgroup of $\mathrm{SL}_{n}\mathbb{F}$ containing only elliptic elements is necessarily bounded [Par03, Théorème 1]. This doesn’t rule out the existence of interesting totally elliptic surface group representations into $\mathrm{SL}_{n}\mathbb{F}$.
1.4. Acknowledgments

### v2

@@ Abstract.
A surface group representation into a Lie group is called totally elliptic if every simple closed curve on the surface is mapped to an elliptic element of the target group. In this note, we characterize all totally elliptic surface group representations into $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ and $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$ by showing that they are either representations into a compact subgroup or Deroin–Tholozan representations.
@@ 1. Introduction
1.1. Main results
Surface group representations are homomorphisms from the fundamental group of a surface $\Sigma$ into a Lie group $G$. The topological quotient of the space of all homomorphisms $\pi_{1}\Sigma\to G$ by the conjugation action of $G$ will be called the character variety of the pair $(\Sigma,G)$ (even though the quotient may not be a Hausdorff topological space) and denoted by
$\operatorname{Rep}(\Sigma,G)=\operatorname{Hom}(\pi_{1}\Sigma,G)/G.$
A character variety is not necessarily a connected space; dynamical behaviour and qualitative properties of representations may vary considerably between components. At one end of the spectrum are components of discrete and faithful representations (so-called higher Teichmüller spaces [Wie18]). This article aims at drawing the reader’s attention towards the other end of the spectrum where connected components are mostly made of non-discrete representations.
We will focus on specific prototype of surface group representations—known as totally elliptic representations—that are antipodal to discrete and faithful ones.
@@ 1.1. Main results
A surface group representation $\pi_{1}\Sigma\to G$ is totally elliptic if every simple closed curve on $\Sigma$ is mapped to an elliptic element of $G$. By an elliptic element of $G$, we mean an element of a compact subgroup of $G$ (sometimes called a compact element of $G$).
The surface $\Sigma$ will always be assumed to be oriented and connected, and may be punctured. A simple closed curve on $\Sigma$ is the free isotopy class of a closed loop on $\Sigma$ that does not self-intersect and is homotopically non-trivial. A simple closed curve can be homotopic to a puncture, in which case we call it a peripheral curve. By definition of a totally elliptic representation, the images of all peripheral curves are elliptic. We want to avoid the situation where peripheral curves are mapped to the identity element in $G$ and we say that a representation is reduced when this is not the case (Definition 2.7).
The first examples of totally elliptic representations, though not the most striking, are those whose image lies in a compact subgroup of $G$. Interestingly, however, by requiring only simple closed curves to be mapped to elliptic elements (as opposed to all curves), one can construct totally elliptic representations with dense image into non-compact Lie groups and whose deformations remain totally elliptic. This turns out to be possible whenever $G$ is a Hermitian Lie group for instance (Section 1.4).
This paper focuses on the case where $G$ is $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ or $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$. The first totally elliptic representations with (Zariski) dense image in $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ were discovered by Benedetto–Goldman [BG99] for 4-punctured spheres, and later generalized by Deroin–Tholozan [DT19] for spheres with arbitrarily many punctures. Those representations were originally named supra-maximal in [DT19], but we shall call them as DT representations (Definition 2.10), to avoid any confusion with maximal representations from higher Teichmüller theory. Deroin–Tholozan proved that deforming conjugacy classes of DT representations, while preserving the rotation angle (Section 2.2) of the peripheral monodromies, leads to compact components inside the corresponding relative character varieties (Section 2.3) which are made exclusively of totally elliptic representations (Theorem 2.9). The question was raised whether DT representations are the only totally elliptic representations into $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, besides representations into compact subgroups which we call orthogonal representations. We bring a positive answer to that question.
@@
Consider a surface $\Sigma$ of genus $g\geq 0$ with $n\geq 0$ punctures, and let $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ be a reduced totally elliptic representation. If $\rho$ is not orthogonal, then $\Sigma$ must be a sphere with at least three punctures ($g=0$, $n\geq 3$), and $\rho$ is necessarily a DT representation.
We prove Theorem A in Section 2. The proof can be sketched as follows. First, we notice that surfaces with genus at least $1$ and different than the closed torus admit simple closed curves that can be represented by the commutator of two fundamental group elements, each representing a simple closed curve (Lemma 2.2). However, the commutator of two elliptic elements of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ is elliptic if and only if it is trivial (Lemma 2.1). These two observations will obstruct the existence of non-orthogonal totally elliptic representations when $g\geq 1$. The more challenging case is when $g=0$. When $\Sigma$ is a punctured sphere, we will prove that totally elliptic and non-orthogonal representations are DT representations by inducting on the number $n$ of punctures on $\Sigma$. The base case $n=4$ is a consequence of a combination of results by Deroin–Tholozan (Proposition 2.12) and Cantat–Loray (Corollary 2.15). For the induction step, we will require the notion of triangle chains developed in [Mar24] and the associated characterization of DT representations (Proposition 2.17).
1.2. Applications and generalizations
An immediate consequence of Theorem A is that DT representations can be characterized inside their relative character variety by the property of being totally elliptic.
Corollary B.
If $\Sigma$ is a sphere with $n\geq 3$ punctures and $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ maps every peripheral curve on $\Sigma$ to a non-trivial elliptic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, then
[…]
The direct implication in Corollary B was established in [DT19, Lemma 3.2] (see also [Mar22, Proposition 2.7]) and the return implication follows immediately from Theorem A. The statement of Corollary B can be refined depending on the peripheral data of $\rho$. The angle of rotation of the image under $\rho$ of a peripheral curve on $\Sigma$ is a called a peripheral angle (Section 2.3). A peripheral angle is always a number in $(0,2\pi)$. When the sum of all peripheral angles of $\rho$ (one for each puncture of $\Sigma$) is not an integer multiple of $2\pi$, then $\rho$ cannot be orthogonal (see Corollary 2.11 for more details). In that case, the statement of Corollary B can be refined to:
If the sum of the peripheral angles of $\rho$ is not an integer multiple of $2\pi$, then
$\rho\text{ is a DT representation}\quad\Longleftrightarrow\quad\rho\text{ is totally elliptic.}$.
In Section 3, we will generalize the statement of Theorem A to surface group representations into $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$—the complexification of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$. Maximal compact subgroups of $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$ are conjugate to the projective unitary group $\operatorname{PSU}(2)$ and surface group representations into a compact subgroup of $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$ are accordingly called unitary representations. In order to generalize Theorem A from $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ to $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$, we need to distinguish between irreducible and reducible representations. Recall that a representation $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$ is reducible if its image is conjugate to a subgroup made of upper triangular elements of $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$; it is irreducible otherwise. Despite the unfortunate linguistic resemblance, the notion of reducible representations and that of reduced representations (Definition 2.7) are different.
Theorem C (Propositions 3.13.2).
Consider a surface $\Sigma$ of genus $g\geq 0$ and with $n\geq 0$ punctures, and let $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$ be a reduced totally elliptic representation.
  • If $\rho$ is irreducible, then either $\rho$ is unitary or $\Sigma$ must be sphere with at least three punctures ($g=0$, $n\geq 3$), and $\rho$ is conjugate to the composition of a DT representation with the inclusion $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}\subset\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$.
  • If $\rho$ is reducible and $g\geq 1$, then the image of $\rho$ is conjugate to a subgroup made of diagonal elements in $\operatorname{PSU}(2)$.
Theorem C leaves open the case of reducible totally elliptic representations into $\mathrm{PSL}_{2}\operatorname{\mathbb{C}}$ for genus-0 surfaces. As it turns out, such representations are not necessarily unitary.
Theorem D (Example 3.4).
If $\Sigma$ is a sphere with at least three punctures ($g=0$, $n\geq 3$), then there exist reduced totally elliptic representations $\pi_{1}\Sigma\to\mathrm{PSL}_{2}\operatorname{\mathbb{C}}$ that are reducible but not unitary.
@@ 1.3. Motivations
Characterizing a certain type of surface group representations by the images of simple closed curves, as we did in Corollary B, may seem baroque at first. Similar problems, however, have sparked substantial interest in recent decades. Bowditch famously raised the question of whether discrete and faithful representations of a closed surface group into $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$—also known as Fuchsian representations—can be characterized by the property of mapping every simple closed curve to a hyperbolic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ (an element whose trace is larger than two in absolute value). In analogy to Definition 1.1, such representations will be called totally hyperbolic.
Question 1 ([Bow98, Question C]).
Let $\Sigma$ denote a closed surface of genus at least two. Is the following equivalence true for a representation $\rho\colon\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$:
$\rho\text{ is a Fuchsian representation}\quad\Longleftrightarrow\quad\rho\text{ is totally hyperbolic}\quad?$
Bowditch’s question (Question 1) has been answered positively for surfaces of genus two by Marché–Wolff [MW16, MW19] and remains open in higher genera. A positive answer to Bowditch’s question would be a finer characterization of Fuchsian representations which are already known to be exactly those representations mapping every non-trivial element of $\pi_{1}\Sigma$ to a hyperbolic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$. A surface group representation into $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ whose image consists of hyperbolic elements and the identity is sometimes called purely hyperbolic, so that Fuchsian representations corresponds to faithful purely hyperbolic representations.
@@
Of course, not every representation is Fuchsian. Goldman proved that the character variety of representations $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, where $\Sigma$ is a closed surface of genus $g\geq 2$, has $34g-$ 3 connected components [Gol88]. The components are distinguished by an integer-valued invariant called Euler number. The two components with extremal Euler number are made of conjugacy classes of Fuchsian representations and form two copies of the Teichmüller space of $\Sigma$; the other $54g-$ 5 components are known as intermediate components and they do not contain any Fuchsian representation. Bowditch’s question (Question 1) is equivalent to asking whether every representation whose conjugacy class lies in an intermediate component maps at least one simple closed curve on $\Sigma$ to a non-hyperbolic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$. Marché–Wolff also proved that a positive answer to Bowditch’s question would imply that the mapping class group action on every intermediate component is ergodic [MW16, Theorem 1.6]. The latter statement is known as Goldman’s ergodicity conjecture [Gol06, Conjecture 3.1] and is open, like Bowditch’s question, for surfaces of genus at least three.
@@
Goldman also conjectured that almost every non-Fuchsian closed surface group representation $\pi_{1}\Sigma\to\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ should be the holonomy of a branched hyperbolic structure on $\Sigma$ which is the data of a hyperbolic metric of $\Sigma$ with several conical singularities of angle $2\pi k$ with $k\geq 2$. To be precise, according to Goldman, every dense representation whose conjugacy class lies in an intermediate component with non-zero Euler number should be the holonomy of a branched hyperbolic structure [Gol06, Conjecture 3.9]. Mathews proved some partial implication from a positive answer to Bowditch’s question to Goldman’s geometrization conjecture for the intermediate component of Euler number $32g-$ 3 [Mat12] and Faraco covered the case of purely hyperbolic representations [Far21].
Bowditch’s question (Question 1) and Goldman’s conjectures on intermediate components are intertwined. In addition to Marché–Wolff’s results, we point out that Goldman’s ergodicity conjecture implies a “full measure version” of the geometrization conjecture, as well as a positive answer to Bowditch’s question on a full measure subset of intermediate components.
@@
Goldman’s two conjectures and Bowditch’s question, as well as the interplay among them, constitute a whole program to study intermediate components which we refer to as the Bowditch–Goldman program.
@@
In this context, totally elliptic representations—particularly DT representations—align naturally with the broader philosophy of the Bowditch–Goldman program, even though, strictly speaking, the program concerns closed surfaces. By definition, they map every simple closed curve to an elliptic element of $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$, and therefore satisfy the analogue of Bowditch’s condition on intermediate representations in a strong sense. The ergodicity of the pure mapping class group action on the character variety of DT representations has been established in [Mar22]. As explained in [DT19, Section 4] and further detailed in [FM23], DT representations are also holonomies of hyperbolic cone structures with branch point singularities, thereby satisfying a precise version of the geometrization conjecture.
The study of totally elliptic surface group representations is also motivated by the recent focus on finite mapping class group orbits in character varieties (see for instance [LLL23, BM24, LT14] and references therein) which led to a complete understanding of finite orbits for surface group representations into $\mathrm{SL}_{2}\operatorname{\mathbb{C}}$. Although not all finite orbits of conjugacy classes of surface group representations into $\mathrm{SL}_{2}\operatorname{\mathbb{C}}$ are made of totally elliptic representations, most of them are (see [BM24] for more details). Totally elliptic representations can thus be viewed as a natural generalization of representations whose conjugacy classes lie in finite mapping class group orbits.
1.4. Perspectives away from $\operatorname{PSL}_{2}\operatorname{\mathbb{C}}$
@@
This note also provides an opportunity to survey recent developments concerning totally elliptic representations into Lie groups $G$ other than $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$. Several families of connected components in relative character varieties—all consisting entirely of conjugacy classes of totally elliptic representations—have been identified in the case where $G$ is a Hermitian Lie group. These components are all compact and arise specifically for punctured spheres. The first examples were found by Tholozan–Toulisse when $G$ is the Lie group $\mathrm{SU}(p,q)$, as well as when $G$ is $\mathrm{Sp}(2n,\operatorname{\mathbb{R}})$ or $\mathrm{SO}^{\ast}(2n)$ [TT21]. Analogous compact components have been identified by Feng–Zhang when $G$ is the identity component of $\mathrm{SO}(2,n)$ [FZ23]. Both group of authors studied the topology of relative character varieties via the non-abelian Hodge correspondence and identified compact components in this way. Not only did they prove that every representation $\rho\colon\pi_{1}\Sigma\to G$ whose conjugacy class belongs to one of the compact components they found is totally elliptic, but they also proved that the local system associated to $\rho$ supports a universal variation of Hodge structure. This means that for every punctured Riemann sphere structure on $\Sigma$, there exists a $\rho$-equivariant holomorphic map from the universal cover $\widetilde{\Sigma}$ to the symmetric space $G/K$. The same property for DT representations was established by Mondello [Mon16].
@@
In contrast to DT representations, which all have Zariski dense image, it remains unclear whether every compact component described in [TT21, FZ23] contains a Zariski dense representation (though almost all of them do, given the prevalence of Zariski dense representations in character varieties). Moreover, the topology of these compact components is still not fully understood, whereas DT components are known to be homeomorphic, and even symplectomorphic, to complex projective spaces [DT19, Theorem 4].
@@
There are however more explicit examples of components in relative character varieties made of totally elliptic representations that were discovered by Marc-Zwecker [MZ24]. Marc-Zwecker’s components are also compact and made of representations of a pair of pants into $\mathrm{SU}(2,1)$ where the peripheral curves are mapped to products of reflections through three complex geodesic lines in the complex hyperbolic space of real dimension four. The three complex geodesic lines are the mirrors and they are arranged as a triangle. The resulting components are homeomorphic to 2-dimensional spheres. The geometric nature of the representations in Marc-Zwecker’s components may lead to other components of totally elliptic representations for spheres with a higher number of punctures obtained by “chaining” the triangles of mirrors in the spirit of [Mar24].
@@
A natural question to ask is whether the class of Lie groups for which there exist (non-trivial) connected components entirely made of totally elliptic surface group representations is larger than the class of Hermitian Lie groups. It is nevertheless unclear whether such components are necessarily compact when the target group is not $\operatorname{PSL}_{2}\operatorname{\mathbb{R}}$ (compare Proposition 2.12), but it seems likely to.
Question 2.
@@
Are there semisimple real Lie groups other than Hermitian Lie groups for which there exist compact connected components of maximal expected dimension in relative character varieties which are made of totally elliptic representations and contain Zariski dense representations?
@@
One should nevertheless be aware that “large” Lie groups, unlike $\mathrm{SL}_{2}\operatorname{\mathbb{R}}$ or $\mathrm{SL}_{2}\operatorname{\mathbb{C}}$, may admit a subgroup $\Gamma$ containing only elliptic elements—a purely elliptic subgroup—but such that $\Gamma$ is not contained in a compact subgroup. [^1] A proof that every purely elliptic subgroup of $\mathrm{SL}_{2}\operatorname{\mathbb{C}}$ is conjugate to a subgroup of $\mathrm{SU}(2)$ can be found in [JS87, Theorem 2.13.1]. The analogous statement for $\mathrm{SL}_{2}\operatorname{\mathbb{R}}$ and $\mathrm{SO}(2)$ is a direct consequence of Lemma 2.1. Bass gives an example of such a $\Gamma$ which is isomorphic to a rank-2 free subgroup of $\mathrm{SL}_{3}\operatorname{\mathbb{C}}$ [Bas80, Counterexample 1.10]. Bass’ example is not Zariski dense and, in fact, it could not be as explained for instance in [Pra94]. However, when $\operatorname{\mathbb{C}}$ is replaced by a non-Archimedean valued field $\mathbb{F}$, then the situation is similar to the case of $\mathrm{SL}_{2}\operatorname{\mathbb{C}}$ as Parreau proved that a finitely generated subgroup of $\mathrm{SL}_{n}\mathbb{F}$ containing only elliptic elements is necessarily bounded [Par03, Théorème 1]. This, of course, does not rule out the existence of totally elliptic surface group representations into $\mathrm{SL}_{n}\mathbb{F}$ whose image is not bounded.
1.5. Acknowledgements